# LayerIF: Estimating Layer Quality for Large Language Models using Influence Functions

### v1

@@ Appendix B Broader Impact
Appendix C Gemma-7b Expert Allocation Results
@@
Table 4: Accuracy accross datasets using AlphaLora, MoLA(3,5,7,8) and LayerIF (+ve) for 160 experts on Gemma-7b.
@@
Appendix D Gemma-7b Sparsity Allocation Results
@@
Appendix E Sparsity Allocation Results with and without smoothing for Mistral-7b-v0.1
@@
We used a window-length of 7 and a polyorder of 3 for our filter. We present the percentage change [^1] Percentage change is computed as (with smoothingwithout smoothingwithout $\left(\frac{\text{with smoothing}-\text{without smoothing}}{\text{without smoothing}}\right)\times 100\%$ in performance at 70% sparsity, comparing results with and without smoothing, for Mistral-7B-v0.1 in Table 6.
@@
Appendix F Statistical Significance of Results
@@
Table 7: Computing the t-test metric to compare LayerIF with Mola and AlphaLora on zero-shot accuracy across the different number of experts (** indicates statistical significance i.e., $p$-value $<0.05$)
@@
Table 8: Computing the t-test metric to compare LayerIF with Alphapruning, Uniform and OWL on zero-shot accuracy across the different number of pruning ratios (** indicates statistical significance i.e., $p$-value $<0.05$)
@@
Appendix G Code and Reproducibility
We follow the datasets and evaluation protocol of AlphaLora [46] for laywer-wise expert allocation and the datasets and evaluation protocol of Alphapruning [19] for layer-wise sparsity allocation. Our code has been anonymized and uploaded here: https://anonymous.4open.science/r/Expert_Allocation-584A and https://anonymous.4open.science/r/LayerIF_Pruning_New-FF34.

### v3

@@ 6 Conclusion
Acknowledgment
We appreciate the reviewers for their insightful comments and suggestions. Hadi Askari and Muhao Chen were supported by the NSF of the United States Grants ITE 2333736, OAC 2531126, and the Amazon Nova Trusted AI Prize. Anshuman Chhabra was supported by the USF Faculty Startup Fund.
@@ Appendix B Broader Impact
Appendix C Clarifying Walk-Through Example
Here, we provide a concrete step-by-step walk-through of our method for greater clarity of our method.
Concrete walk-through for expert allocation (CommonQ dataset, 32-layer LLM, T=160 experts, $\beta=3$)
Below we provide the exact numbers produced by our implementation. For brevity only the first 5 layers are shown; the remaining layers follow identically.
Layer $i$ Raw IF score $v_{i}$ Inverted $\tilde{v}_{i}$ Power-scaled $\hat{v}_{i}$ Fractional $f_{i}$ Final experts $s_{i}$
0 $5.01\times 10^{11}$ $5.01\times 10^{11}$ $6\times 10^{35} 1.12 2$
1 $8.36\times 10^{11}$ $8.36\times 10^{11}$ $5.84\times 10^{35}$ 5.21 6
2 $7.99\times 10^{11}$ $7.99\times 10^{11}$ $5.10\times 10^{35}$ 4.55 6
3 $8.46\times 10^{11}$ $8.46\times 10^{11}$ $6.05\times 10^{35}$ 5.40 6
4 $8.16\times 10^{11}$ $8.16\times 10^{11}$ $5.43\times 10^{35}$ 4.84 6
$\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$
  • Step1 (sign inversion). Invert the negative values to positive.
  • Step2 (power transform). Amplify disparities by raising to the power $\beta$ to get $\hat{v}$.
  • Step3 (normalise to budget). Scale $\hat{v}$ so that the fractional allocations $\sum f_{i}$ equal $\mathbf{T-m=128}$ extra experts ($m=32$ layers).
  • Step4 (floor + redistribute). Set $s_{i}=\lfloor f_{i}\rfloor+1$, guaranteeing $\geq 1$ expert/layer, then distribute the remaining 24 units to the layers with the largest fractional parts.
Finally, the complete allocation for CommonQ becomes:
[2,6,6,6,6,5,7,6,8,7,6,7,6,6,6,5,6,6,5,4,5,5,4,2,4,5,3,4,2,3,4,3][2,6,6,6,6,5,7,6,8,7,6,7,6,6,6,5,6,6,5,4,5,5,4,2,4,5,3,4,2,3,4,3]
Interpretation. The middle layers receive the most experts, indicating they are comparatively under-trained for CommonQ; the first and very deep layers receive lower number of experts, reflecting higher estimated quality.
Appendix D Gemma-7b Expert Allocation Results
@@
Table 5: Accuracy across datasets using AlphaLora, MoLA(3,5,7,8) and LayerIF (+ve) for 160 experts on Gemma-7b.
@@
Appendix E Gemma-7b Sparsity Allocation Results
@@
Appendix F Sparsity Allocation Results with and without smoothing for Mistral-7b-v0.1
@@
We used a window-length of 7 and a polyorder of 3 for our filter. We present the percentage change [^3] Percentage change is computed as $\left(\frac{\text{with smoothing}-\text{without smoothing}}{\text{without smoothing}}\right)\times 100\%$ in performance at 70% sparsity, comparing results with and without smoothing, for Mistral-7B-v0.1 in Table 7.
@@
Appendix G Statistical Significance of Results
@@
Table 8: Computing the t-test metric to compare LayerIF with Mola and AlphaLora on zero-shot accuracy across the different number of experts (** indicates statistical significance i.e., $p$-value <0.05<0.05)
@@
Table 9: Computing the t-test metric to compare LayerIF with Alphapruning, Uniform and OWL on zero-shot accuracy across the different number of pruning ratios (** indicates statistical significance i.e., $p$-value <0.05<0.05)
@@
Appendix H Wall Clock Times and Memory Consumption for Varying Dataset and Model Sizes
We report the wall-clock times of our Influence Function (IF) computation using the DataInf approximation, compared against the TracIn baseline, under varying dataset sizes and model scales. All experiments were performed on $4\times$ NVIDIA RTX 6000 Ada Generation GPUs.
H.1 Scalability to Dataset Size
Table 10 presents runtimes for computing IFs over a single layer with increasing numbers of training and validation samples. As expected, the runtime scales approximately linearly with dataset size. Although DataInf is substantially slower than TracIn, the total cost remains modest relative to the computational expense of pretraining large language models. For comparison, training LLaMA2-70B reportedly required $\sim$1.72M GPU hours, whereas DataInf adds only a minor diagnostic overhead that can be amortized across layers to identify underperforming components during training.
Table 10: Wall-clock runtimes (in seconds) for computing IFs for a single layer under different dataset sizes on Mistral-7B-v0.1.
Dataset Size (Train×Val) Method Runtime (s)
3000×500 DataInf 163,721
TracIn $1.35\times 10^{-2}$
300×50 DataInf 402.77
TracIn $3.31\times 10^{-5}$
30×5 DataInf 15.3
TracIn $1.19\times 10^{-6}$
The peak GPU memory consumption during influence computation for a dataset of 300 training and 50 validation samples was approximately 35000MiB for the 7B model. This footprint remains well within the typical per-GPU memory budget, indicating that influence-based diagnostics can be feasibly integrated into existing training pipelines.
H.2 Scalability to Model Size
We also evaluate scaling across model sizes in Table 11. The runtime grows roughly proportionally to parameter count.
Table 11: Wall-clock runtimes (in seconds) for DataInf and TracIn under different model sizes (300×50 dataset).
Model Method Runtime (s)
Qwen2.5-3B DataInf 58.063
TracIn $2.79\times 10^{-5}$
Qwen2.5-32B DataInf 2,824.16
TracIn $6.82\times 10^{-5}$
Overall, while DataInf incurs higher per-layer computation cost than first-order approximations such as TracIn, it remains computationally tractable even for multi-billion-parameter models. Its cost is negligible relative to the full pretraining budget, while providing rich layer-wise diagnostic signals that can inform targeted retraining or pruning strategies.
Appendix I Code and Reproducibility
We follow the datasets and evaluation protocol of AlphaLora [47] for layer-wise expert allocation and the datasets and evaluation protocol of Alphapruning [19] for layer-wise sparsity allocation. Our code has been anonymized and uploaded here: https://github.com/HadiAskari/Expert_Allocation and https://github.com/HadiAskari/LayerIF_Pruning_New.